OpenStudy (h0pe):

When N is divided by 10, the remainder is a. When N is divided by 13, the remainder is b. What is N modulo 130, in terms of a and b? (I need my answer in the form ra+sb, where r and s are replaced by nonnegative integers less than 130.)

OpenStudy (rational):

familiar with congruences ?

OpenStudy (h0pe):

Somewhat

OpenStudy (rational):

somewhat is enough :)

OpenStudy (rational):

We're given \[N\equiv a\pmod{10}\\N\equiv b\pmod{13}\] second congruence is same as \(N=13k+b\tag{1}\) substitute that in first congruence and solve \(k\)

OpenStudy (h0pe):

So that gives me \[10x+a=13k+b\]?

OpenStudy (h0pe):

How do I solve it from there?

OpenStudy (rational):

yes, your goal is to solve \(k\) \(10x+a=13k+b\) \(10x + a = 10k+3k+b\) \(10(x-k)+a-b = 3k\) \(10y+a-b = 3k\) fine so far ?

OpenStudy (h0pe):

Yes, but how am I supposed to solve it from there?

OpenStudy (rational):

multiply both sides by "7"

OpenStudy (h0pe):

Why?

OpenStudy (rational):

\(10y+a-b=3k\) multiply both sides by 7 and get \(10z+7(a-b) = 20k+k\) \(10w+7(a-b)=k\)

OpenStudy (rational):

substitute that in equation \((1)\) and you're done!

OpenStudy (h0pe):

Is \[N=13k+b\] equation 1?

OpenStudy (rational):

Yep

OpenStudy (h0pe):

So \[13(10w+7(a-b))\]?

OpenStudy (h0pe):

well \[N=13(102+7(a-b))+b\]?

OpenStudy (h0pe):

10w I menat

OpenStudy (h0pe):

gah so many typos today \[N=13(10w+7(a-b))+b\]

OpenStudy (rational):

\[\begin{align}N &= 13k+b\\~\\ &=13[10w+7(a-b)] +b\\~\\ &=130w+\color{blue}{91(a-b)+b} \end{align}\]

OpenStudy (h0pe):

and to put it in the form of \[ra+sb\]? \[130w+91a-90b\]

OpenStudy (h0pe):

How do I put that into the ra+sb form?

OpenStudy (rational):

91a - 90b is the remainder when u divide N by 130

OpenStudy (rational):

r = 91 s = -90

OpenStudy (h0pe):

Oh... thank you!

OpenStudy (rational):

ofcourse if u want positive numbers -90 is same as 40 in mod 130

OpenStudy (rational):

r = 91 s = 40 looks more better

OpenStudy (h0pe):

yeah :)